OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748. -172,186,884 -57,395,628 57,395,628 172,186,884

OpenStudy (cwrw238):

common ratio 7th root of -8748 / 4 = -3 16th root = 4* -3^15 = -57,395,628

OpenStudy (cwrw238):

its 7th root because a8 / a1 = ar^7 / a = r^7

OpenStudy (cwrw238):

*a1 r^7 / a1

OpenStudy (anonymous):

\[a_n = a_1*r^{n-1} \]\[a_8 = a_1*r^7 \]\[-8,748 = 4*r^7\]\[r^7 = -2187\]\[r=-3\]

OpenStudy (anonymous):

\[a_{16}=4*(-3)^{15}\]\[a_{16}=57,395,628\]

OpenStudy (anonymous):

Ok this one's correct. Sorry..

OpenStudy (anonymous):

haha thank you! i was like wait a minute lol

OpenStudy (anonymous):

agh it's negative! -57,395,628

OpenStudy (anonymous):

Something's wrong with my keyboard.. :/

OpenStudy (anonymous):

thanks! :]